# More than Meets the Eye: Understanding the Effect of Individual Objects on Perceived Visual Privacy

### v1

@@ Abstract.
Visual Privacy, Survey, Social Media/Online Communities, Empirical Studies
copyright: none
1. Introduction
00footnotetext: https://cocodataset.org/#explore; url to image:
http://farm3.staticflickr.com/2596/3855197434_72ccaa2ed3_z.jpg
With the advent and ubiquity of affordable camera technologies, capturing and sharing visual content has become the new norm of everyday social interactions. A study in 2025 (Broz, 2023) reports that 14 billion images are shared daily on social media networks and instant messengers, a high percentage of which include photos of people with friends and family members. However, by sharing photos of their private lives and personal moments, they may inadvertently jeopardize their privacy. Exposing personally identifiable or sensitive information online can have serious consequences, such as identity theft (Burnes et al., 2020), cyber bullying (Chan et al., 2019), or stalking (Kaur et al., 2021).
One of the well-studied privacy concerns related to sharing content online is behavioral profiling and targeted advertisements. These systems mainly rely on users’ search queries, browsing patterns, and metadata to transform personal data and privacy into a commodity. However, what is less known, at least to the general public, is the significant amount of information that can be harvested purely from the visual content by human adversaries and automated systems. Figure 4 demonstrates an example of how much information a single visual content can reveal by combining different visual cues extracted from the image. Investigative journalism led by media outlets  [^1] The New York Times visual investigations — https://www.nytimes.com/spotlight/visual-investigations and advocacy groups  [^2] Bellingcat investigations — https://www.bellingcat.com/category/news/?fwp_categories=news have demonstrated how open-source intelligence (OSINT) tools and techniques can be utilized to extract rich information from shared visual content that was not initially intended (Mossou, 2023; Weide, 2023; Urbina, 2020). Similarly, accessibility-focused HCI research works have also explored privacy concerns associated with intentional or accidental disclosure of visual information (Akter et al., 2020b, a; Stangl et al., 2020; Zhang et al., 2024). However, there is still some uncertainty about whether and how users can understand the complex visual and contextual cues in the photos that they share online. We explore this research theme in our work to uncover the nuances of perceived visual privacy.
Our goal is to understand how individual objects (i.e., the visual elements within the image), which may have varying privacy risks in different contexts, influence users’ privacy decisions. Previous studies focused mainly on the perceived privacy of the entire image to understand users’ perceptions and behaviours while sharing photos online. More specifically, these works have focused on understanding whether a user evaluates the whole image as private or not  (Patwari et al., 2024; Tonge and Caragea, 2020), or focuses on assessing the privacy risks of a single object (Orekondy et al., 2017). However, these approaches overlook the effect of context and the co-presence of different privacy-sensitive objects within an image. A few works on users’ privacy preferences regarding the sharing of visual content on social media have examined spatial context (location where the photo was taken) and social context (e.g., the number of individuals present in the image and the subject’s relation to the image) (Hoyle et al., 2020). Our work extends this line of research by exploring more fine-grained aspects of the visual content.
@@ 1. Introduction
Figure 1. An example demonstrating the amount of information revealed (intentionally or inadvertently) through a single visual content. The image is sampled from the MS-COCO dataset explorer44footnotemark: 4and the text descriptions are a combined interpretation by a human observer (one of the authors of this work) and commercial GenAI methods.
@@
Contributions: Our main contributions from this work are as follows. First, we offer novel insights into how end-users evaluate visual privacy and the factors that shape their privacy perceptions. Secondly, our work provides empirical evidence on users’ privacy heuristics, derived from a mixed-methods analysis of qualitative and quantitative data gathered through an online user survey (n = 92). Finally, we demonstrate the feasibility of using synthetic images in user studies, highlighting their value as an alternative when obtaining real-world data is challenging.
@@
Overview of results: Our results are drawn by fixing the individual visual element (referred to as objects) in the foreground and studying its perceived privacy by the user using varying combinations of objects in the background. In such a setup, we make the following observations.
@@
  • We found that end-users demonstrate a latent cognitive model for evaluating the intricate visual and contextual cues of the images they publish online. Users pay attention to granular details about the objects, such as what or who is in the photo and the situational context in which the photo is captured. This finding supports the well-established theories of contextual anchors of privacy decisions in the prior art (Nissenbaum, 2004), but also extends the knowledge by providing evidence about micro-details of context.
@@
  • We found that the presence of specific categories of sensitive objects in the background dictates the privacy perception of those in the foreground. Especially when the background objects are unrelated to each other, the presence of other objects has a negligible impact on the users’ foreground privacy perception due to the dominance effect. While previous studies have discussed the impact of similar salient features on privacy perception using anecdotal examples, we provide empirical evidence for the dominance effect.
@@
  • Certain combinations of semantically related sensitive objects in the background also contribute to the perceived privacy of those in the foreground. While such individual background objects may still influence users’ privacy judgments to some extent, their co-presence can trigger stronger concerns than what each would cause in isolation. Prior research has extensively examined inference risks from the combination of related elements from different datasets. However, a similar effect emerging from a single data item (such as individual images) is less explored to the best of our knowledge. In this direction, our findings offer novel insights into pair-wise co-presence effect and its potential to trigger privacy concerns.
@@
Our work contributes towards a comprehensive understanding of human visual privacy perception that can potentially be utilized for developing personalized privacy preference options for sharing visual content online. In particular, the insights from this study can be leveraged for designing assistive technologies that share the entire content while removing, blurring, or masking privacy-sensitive objects, but still function effectively. Such technologies can be integrated into various AI applications, including privacy-sensitive visual aid devices, smart home assistants, wearable cameras, augmented reality platforms, surveillance systems, and live classroom environments.
2. Background and Related Work
2.1. Privacy Risks in Visual Content
Online platforms increasingly rely on AI-driven personalization and pervasive data collection practices that amplify privacy risks. Through everyday interactions, and often without full awareness, users reveal personal and behavioral data that can be profiled to infer sensitive attributes (Goyeneche et al., 2023; Shrestha et al., 2024) such as health status, political beliefs, or identity. Visual media, such as photos and videos, are particularly information-rich, with shared content potentially exposing location, social ties, or sensitive activities, leaving users vulnerable to reputational damage, discrimination, or harassment. Restricting privacy settings does not fully protect users, as both the platform and third-party algorithms can still access and repurpose these images in ways beyond the user’s intent, thereby increasing privacy risks (Liu et al., 2022; Amil et al., 2024). Certain user groups, such as children and people with visual impairments, face unique challenges in recognizing and controlling these risks (Stangl et al., 2022; Akter et al., 2020b; Goyeneche et al., 2023). Moreover, advances in computer vision amplify these risks, as increasingly powerful models can analyze and recognize visual content with high accuracy, enabling new forms of information that adversaries can easily exploit (Zhao et al., 2025).
2.2. Perceived Privacy, Norms and Behavior
Perceptions of visual privacy are shaped not only by what is visible in an image, but also by expectations of who will see it, how the platform mediates sharing, and the broader social norms surrounding disclosure. Contextual integrity theory (Nissenbaum, 2004) also supports this: privacy is maintained when information flows conform to the norms and expectations of a given context. Empirical work demonstrates that situational and normative factors, such as intention, relationships, or setting of sharing, often matter more than the presence of objects in an image (Hoyle et al., 2020). Building on this perspective, previous research has also applied contextual properties to guide the design of privacy protections for online visual content (Liu et al., 2022). Our work extends these insights by examining how object-level cues and their co-occurrence interact with contextual factors to influence perceived privacy.
2.3. Visual Privacy Protection
To mitigate the risks mentioned above, prior work has explored privacy-enhancing technologies for visual content. Technical solutions include automatic detection and obfuscation of sensitive regions, blurring or masking private content, facial de-identification, redaction before sharing (Ravi et al., 2024), or developing AI-assisted tools that guide users in making informed decisions (Orekondy et al., 2017). Other studies emphasized user-centered controls, such as allowing people to set sharing preferences on upload or delegate privacy decisions to trusted assistants (Stangl et al., 2022). However, many existing tools do not offer robust protection mechanisms or often overlook how specific objects and contextual cues shape perceptions of privacy.
Despite current limitations, visual privacy protection efforts and understanding user privacy practices align with broader regulatory principles, such as the General Data Protection Regulation (GDPR) in the European Union, which treats visual data as personal data, granting individuals rights over how it is collected, processed, and shared. More recently, the EU AI Act has highlighted that AI systems using personal and biometric data (the majority of them are extracted from visual content) should be labeled as high-risk systems, requiring stronger safeguards and oversight. These frameworks highlight the need for privacy-aware systems that can detect, interpret, and protect sensitive visual cues in ways that adhere to legal obligations and respect users’ preferences.
2.4. Preliminaries
In line with previous research (Akter et al., 2020b; Niu et al., 2025; Orekondy et al., 2018), we define the key concepts used in our research and explain the relationships between them.
@@ 2.4. Preliminaries
  • Object denotes to a bounded, semantically meaningful entity that belongs to a specific category and can be distinguished from other objects and the background. In a visual content, visual objects (e.g., face, person, tree) are entities that can be recognized based on their visual appearance such as shape, texture, color, while textual objects (e.g., name, date) consist of symbolic representations (characters, words, phrases).
  • Privacy-sensitive object (PSO) refers to an object that contains personally identifiable or sensitive information that owners may feel uncomfortable sharing on public platforms. A visual content can include PSOs and non-PSOs (objects that do not carry sensitive information) and background. Referring again to Figure 4, person in the center of the photo and the bystanders are labeled PSO, the trees are labeled non-PSO, and the sky and floor belong to the background.
@@
  • Foreground PSO is the primary subject of interest and is typically located in the center of a visual content, while the background PSO contains sensitive information but is not treated as the main subject. In Figure 4, person in the center is the foreground PSO and the bystanders are the background PSO.
@@
  • Perceived privacy level (PPL) is the degree to which individuals believe their personal information, activities, or beliefs are protected from unwanted observation, access, or misuse. It is subjective as the same setting might be judged differently by various individuals.
@@
Building upon this terminology, we construct the following hypotheses:
H1 The perceived privacy level of a foreground PSO increases when it is accompanied by a background PSO.
H2 The surrounding visual scene affects the perceived privacy level of the same PSO present in the image.
H3 The co-presence of multiple background PSOs further elevates the perceived privacy level of a foreground PSO compared to when only a single background PSO is present.
3. Methodology
3.1. Study Design
3.1.1. Selection of Environments
We consider two different environments for the survey: the café and the office. These environments represent distinct contexts, one social and one professional, in which photography is taken and might also be shared. Privacy expectations are usually lower in café due to the co-presence of strangers and a sense of privacy through anonymity. On the other hand, people in office environment has clear roles, often handle professional or sensitive information, leading discomfort to take photos and share in social media.
3.1.2. Selection of Foreground and Background PSOs
To determine both foreground and background PSOs, we examined the subset of the VISPR dataset designed for image redaction (VISPR-Redacted (Orekondy et al., 2018)) and the VizWiz-Priv dataset (Gurari et al., 2019). VISPR-Redacted contains user uploaded, publicly available Flickr images and is specially curated for automatic detection of PSOs and redacting them by masking. VISPR-Redacted contains 24 labels, all of which are considered PSO. VizWiz-Priv is a large-scale collection of real-world images captured by blind photographers and accompanied by questions and crowd-sourced answers. VizWiz-Priv includes 23 categories that annotators label as PSO. Our analysis revealed that both datasets include several identical labels. This provided a perspective on which PSOs are commonly shared with others and online. We chose the three most common object categories from VizWiz-Priv to serve as our foreground PSOs: face, miscellaneous paper and computer screen. Each foreground PSO carry inherent sensitive information while also offering potential information for users:
  1. (1)
@@ 3.1.2. Selection of Foreground and Background PSOs
For each foreground PSO, we identified five background PSOs through an internal expert discussion process on both datasets, considering diversity in modality (visual vs. textual), potential sensitivity, and variation in real-world presence. These background PSOs can change user comfortability and can be present with the associated foreground PSO. Our set of foreground/background PSO combinations is presented in Table 1.
Table 1. List of foreground and background PSOs.
@@
Face face (another, a.), poster, medicine, tattoo, landmark
Miscellaneous paper face (photo, p.), full name, address, birth date, signature
Computer/phone screen (screen) face (reflection, r.), date, e-mail, username, phone number
3.1.3. Image Generation
To ensure that the participants’ judgments were not influenced by uncontrolled factors such as inconsistent lighting, backgrounds, or photographic context, we deliberately avoided real images. Instead, we relied on synthesized images so that visual conditions remained as similar as possible across variations, allowing participants to focus on the intended differences between foreground and background PSOs.
We used OpenAI’s SORA text-to-image model  [^5] https://sora.chatgpt.com/ to generate all images. For images containing only a foreground PSO, we designed comprehensive prompts to produce realistic and contextually meaningful scenes. To maintain visual consistency, we applied SORA’s Remix functionality: instead of regenerating an entire image, we selected specific regions and instructed the model to add the required background PSO within that context. This approach preserved lighting, perspective, and composition in all conditions. We used a male subject for the office environment and a female subject for the café.

### v2

@@ Abstract
I Introduction
@@
Figure 1: An example demonstrating the amount of information revealed (intentionally or inadvertently) through a single visual content. The image is sampled from the MS-COCO dataset explorer and the text descriptions are a combined interpretation by a human observer (one of the authors of this work) and commercial GenAI methods (link to actual image is here).
With the advent and ubiquity of affordable camera technologies, capturing and sharing visual content has become the new norm of everyday social interactions. A study in 2025 [photutorial2023] reports that 14 billion images are shared daily on social media networks and instant messengers, a high percentage of which include photos of people with friends and family members. However, by sharing photos of their private lives and personal moments, they may inadvertently jeopardize their privacy. Exposing personally identifiable or sensitive information online can have serious consequences, such as identity theft [burnes2020risk], cyber bullying [Chan03042019], or stalking [kaur2021systematic].
One of the well-studied privacy concerns related to sharing content online is behavioral profiling and targeted advertisements. These systems mainly rely on users’ search queries, browsing patterns, and metadata to transform personal data and privacy into a commodity [boerman2017online, puglisi2017web]. However, a significant amount of information can be harvested purely from the visual content by humans and computer vision systems. Figure 1 demonstrates an example of how much information a single photo can reveal by combining different visual cues extracted from the image. News media investigations  [^1] The New York Times visual investigations and Bellingcat investigations have demonstrated how open-source intelligence tools and techniques can be utilized to extract rich information from shared visual content that was not intended initially [ww2photomystery2023, chronolocation2023, ChinaShips2020]. Similarly, accessibility-focused research works in the field of human-computer interaction (HCI) have also explored privacy concerns associated with intentional or accidental disclosure of visual information [akter2020uncomfortable, akter2020privacy, stangl2020visual, zhang2024designing]. However, there is still some uncertainty about whether and how users can understand the complex visual and contextual cues in the photos they share online. Our work aims to explore this research theme to uncover the nuances of perceived visual privacy.
Our goal is to understand how individual objects (i.e., the visual elements within the image), which may have varying privacy risks in different contexts, influence users’ privacy decisions. Previous studies focused mainly on the perceived privacy of the entire image to understand users’ perceptions and behaviors while sharing photos online. More specifically, these works have focused on understanding whether a user evaluates the whole image as private or not  [patwari2024Perceptanon, Tonge2020Image, habib2019impact, ahern2007over], or focuses on assessing the privacy risks of a single object [orekondy2017towards]. However, these approaches overlook the effect of context and the co-presence of different privacy-sensitive objects within an image. A few works on users’ privacy preferences regarding the sharing of visual content on social media have examined spatial context (location where the photo was taken) and social context (e.g., the number of individuals present in the image and the subject’s relation to the image) [hoyle2020privacynorms]. Our work extends this line of research by exploring more fine-grained aspects of visual content.
@@
Contributions: We offer novel insights into how end-users evaluate visual privacy and the factors that shape their privacy perceptions. Secondly, we provide empirical evidence on users’ privacy heuristics, derived from a mixed-methods analysis of qualitative and quantitative data gathered through an online user survey (n = 92). Finally, we show the feasibility of using synthetic images in user studies, as a valuable alternative when real-world data is difficult to obtain.
@@
Overview of results: We draw results by fixing the individual visual element (referred to as objects) in the foreground and studying its perceived privacy by the user using varying combinations of objects in the background. In summary, we make the following observations.
@@
  • We found that end-users demonstrate a latent cognitive model for evaluating the intricate visual and contextual cues of the images they publish online. Users pay attention to granular details about the objects, such as what or who is in the photo and the situational context in which the photo is captured. This finding supports the well-established theories of contextual anchors of privacy decisions [nissenbaum2004privacy], and extends the knowledge by providing evidence about micro-details of context.
@@
  • We found that having certain categories of sensitive objects in the background dictates the privacy perception of those in the foreground. Especially when the background objects are unrelated to each other, the presence of other background objects has a negligible impact on the users’ privacy perception of the foreground object due to the dominance effect. While previous studies have discussed the impact of similar salient features on privacy perception using anecdotal examples, we provide empirical evidence for the dominance effect.
@@
  • Certain combinations of semantically related sensitive objects in the background contribute to the perceived privacy of those in the foreground. While such individual background objects may still influence users’ privacy judgments to some extent, their co-presence can trigger stronger concerns than what each would cause in isolation. Prior research has investigated the inference risks that arise when related elements from different datasets. However, a similar effect emerging from a single data item (such as individual images) is less explored to the best of our knowledge. In this direction, our findings offer novel insights into pair-wise co-presence effect and its potential to trigger privacy concerns.
@@
Our work contributes towards a comprehensive understanding of visual privacy perception that can potentially be utilized to develop personalized privacy preference options for sharing visual content online. In particular, the insights from this study can be leveraged for designing assistive technologies that share the entire content while removing, blurring, or masking privacy-sensitive objects, but still function effectively. Such technologies can be integrated into various AI applications, including privacy-sensitive visual aid devices, smart home assistants, wearable cameras, augmented reality platforms, surveillance systems, and live classroom environments.
II Related Work
II-A Contextual Foundations of Visual Privacy
Many privacy theories argue that human perception of privacy is both content- and context-dependent [wisniewski2022privacy]. Rather than treating privacy as a static property of information, these theories highlight how people interpret its meaning, intent, and appropriateness in the moment. For example, contextual integrity theory argues that privacy is about whether information flows align with the social norms of a given context, such as who receives the information, what is being shared and how it is shared [nissenbaum2004privacy]. Similarly, boundary regulation [petronio2002boundaries] and privacy calculus [dinev2006extended, acquisti2007can] treat privacy as a negotiation outcome of interpersonal boundaries with others and of perceived benefits of information disclosure. Empirical HCI research supports these theoretical perspectives, showing that users’ information sharing decisions depend heavily on situational norms and social relationships. Prior work on visual privacy has explored, e.g., how user dispositions (e.g., demographics, personal traits), the aesthetics of the content (e.g., location, people), the intended audience, the presence of bystanders, and the surrounding environment in which a photo is taken can influence privacy perceptions [habib2019impact, ahern2007over, hoyle2020privacynorms, Niu2025Bystanders, kairam2016snap]. Collectively, these findings indicate that visual privacy judgments emerge from contextual signals and micro-details of the content. Our work builds on these foundations and empirical results by examining privacy perception under controlled manipulation of micro-details and contextual conditions. While prior work has focused mainly on whole image, we investigate how object-level cues and their interactions contribute to perceived privacy.
II-B Visual Privacy Risks and Protection
A wide range of privacy risks associated with visual content have been documented in the research literature. Such risks emerge from the distinct and recognizable elements contained in the image, often revealing more information than the sharer intends. Prior work has shown that photos can expose biometric or demographic details, geolocation, social relationships and status, or health and economic conditions [li2020towards, 10.1145/3547299, goyeneche2024linked]. Individuals who did not voluntarily participate in image creation (e.g., bystanders, passersby, or children) could also be exposed when the image is shared without their knowledge or consent [akter2020uncomfortable, akter2020privacy, Niu2025Bystanders, 10.1145/3517384]. Privacy risks are further amplified by sensitive attributes inferred at scale from image content and metadata by modern data-harvesting systems [amil2024impact]. These findings illustrate that privacy risks in images arise from inferences that can be drawn from how visual and contextual cues relate to each other within a scene.
To mitigate the risks mentioned above, prior work has explored privacy-enhancing technologies for visual content [10.1145/3708501, wen2024image, chiang2025understanding]. Technical solutions include automatic detection and obfuscation of sensitive details using blurring, redaction, and facial de-identification [orekondy2017towards, ravi2024review, khamis2022deepfakes, vishwamitra2017blur]. On the other hand, user-centric solutions offer options to restrict the sharing of visual content to a specific set of people or to delegate privacy decisions to trusted assistants [zhang2024designing, hoyle2020privacynorms, 10.1145/3517384, 10.1145/2396761.2398735]. However, these solutions treat all sensitive elements as uniformly risky, overlooking how the co-presence of multiple elements impacts users’ privacy perception. The mismatch between existing protection mechanisms and human perception emphasizes the need for a deeper understanding of object-level interactions, which is the focus of our work.
II-C Visual Privacy Methodological Approaches
Research on visual privacy has traditionally relied on natural images collected from social media, photo-sharing platforms, or curated datasets. Prior studies have used such images in user studies, where participants are asked to report, e.g., their comfort level in sharing them with others, their intended audience, the control mechanisms they use, and the reasoning behind their decisions, which stem from participants’ privacy preferences. Furthermore, many studies have leveraged photos provided by participants to investigate which types of content are considered sensitive. Researchers have also studied how the visibility of certain objects affects privacy judgments by obfuscating sensitive regions of the image before measuring participants’ privacy perceptions [orekondy2017towards, li2017effectiveness, khamis2024perspectives]. Despite ecological validity and high-level insights, the main limitation of using natural images is the lack of control over modulating specific visual aspects. As a result, it is difficult to observe the role of individual foreground or background objects, their co-presence and interactions, or changes in the surrounding environment that shape perceived privacy. To overcome such limitations, synthetic images produced by modern generative models that offer the controlled manipulation of visual features can be a helpful approach. Recent HCI research has explored whether humans can differentiate between synthetic and real faces [lu2023seeing, lago2021more, nightingale2021synthetic, shen2021study]. Neurophysiological studies also demonstrate that synthetic images can approximate human perceptual judgments in controlled settings and motivate the use of synthetic stimuli in experimental research [bilucaglia2024emotional]. Building on this direction, our study employs carefully designed prompts to control the components of images during the image generation process. We use these synthetically generated images as stimuli in a user study aimed at investigating how object-level features and contextual factors shape privacy perception.
III Methodology
III-A Preliminaries
In line with prior work [akter2020uncomfortable, Niu2025Bystanders, orekondy2018connecting], we define the key concepts and their relationships as follows.
@@ III-A Preliminaries
  • Object denotes a bounded, semantically meaningful entity that belongs to a specific category and can be distinguished from other objects and the background. The visual objects (e.g., face, person, tree) are entities that can be recognized based on their visual appearance such as shape, texture, color, while textual objects (e.g., name, date) consist of symbolic representations (characters, words, phrases).
  • Privacy-sensitive object (PSO) refers to an object that contains personally identifiable or sensitive information that owners may feel uncomfortable sharing on public platforms. A visual content generally consists of objects with different levels of sensitivity, along with the background. An object may be classified as a PSO either due to explicitly defined data protection regulations or visual indicators such as the surrounding context. As illustrated in Figure 1, the person in the center of the image and the nearby bystanders are marked as PSO (since they constitute personally identifiable information under GDPR), whereas the sky and the floor belong to background. Other objects, such as trees or trash bins, can also be elevated to PSO when the goal is to infer the geographical location where the image was captured.
@@
  • Foreground PSO is the primary subject of interest and is typically located in the center of visual content, despite the presence of other sensitive elements in the background. In Figure 1, the person in the center is the foreground PSO and the bystanders are the background PSO.
@@
  • Perceived privacy level (PPL) is the degree to which individuals subjectively believe their personal information, activities, or beliefs are protected from unwanted observation, access, or misuse.
@@
Based on this terminology, we construct hypotheses as follows:
H1 : The perceived privacy level of a foreground PSO increases when it is accompanied by a background PSO.
H2 : The surrounding visual scene affects the perceived privacy level of the same PSO present in the image.
H3 : The co-presence of multiple background PSOs further elevates the perceived privacy level of a foreground PSO compared to when only a single background PSO is present.
III-B Study Design
III-B1 Selection of Environments
We consider two survey environments: a café and an office. These represent distinct social and professional contexts, both communal spaces where photos may be taken. Privacy expectations are usually lower in cafés due to the sense of privacy through anonymity. In offices, clear roles and the handling of sensitive or private information can make people uncomfortable with taking and sharing photos on social networks.
III-B2 Foreground and Background PSOs
To determine both foreground and background PSOs, we examined the subset of the VISPR dataset designed for image redaction (VISPR-Redacted [orekondy2018connecting]) and the VizWiz-Priv dataset [Gurari2019VizWiz]. VISPR-Redacted contains user uploaded, publicly available Flickr images and is curated for automatic detection of PSOs and redacting them by masking. VISPR-Redacted contains 24 labels, all marked as PSOs. VizWiz-Priv is a large-scale collection of real-world images captured by blind photographers and accompanied by questions and crowd-sourced answers. VizWiz-Priv includes 23 categories that annotators label as PSO. Our analysis revealed that both datasets include several identical labels, providing insights into PSOs that are commonly shared with others and online. We chose the three most common object categories from VizWiz-Priv as our foreground PSOs: face, miscellaneous paper and computer screen. Each foreground PSO carry inherent sensitive information while also offering potential information for users:
  1. 1.
@@ III-B2 Foreground and Background PSOs
For each foreground PSO, we identified five background PSOs through an internal expert discussion on both datasets, considering diversity in modality (visual vs. textual), potential sensitivity, and their frequency of being together with foreground PSOs. Our set of foreground/background PSO combinations is presented in Table I.
TABLE I: List of Foreground and Background PSOs.
@@
Face Face (another, a.), Poster,
Medicine, Tattoo, Landmark
Miscellaneous Face (photo, p.), Full name,
paper Address, Birth date, Signature
Computer/phone screen Face (reflection, r.), Date,
(screen) E-mail, Username, Phone number
III-B3 Image Generation
To prevent participants’ judgments from being influenced by uncontrolled factors such as inconsistent lighting, backgrounds, or photographic context, we deliberately avoided real images. Instead, we relied on synthesized images to keep visual conditions consistent across variations, allowing participants to focus on the intended differences between foreground and background PSOs.
We used OpenAI’s SORA text-to-image model  [^2] https://sora.chatgpt.com/ to generate all images. For images containing only a foreground PSO, we designed comprehensive prompts to produce realistic and contextually meaningful scenes. To maintain visual consistency, we applied SORA’s Remix functionality: instead of regenerating an entire image, we selected specific regions and instructed the model to add the required background PSO within that context. This approach preserved lighting, perspective, and composition under all conditions. The generated images did not deviate significantly from each other, so the focus remains on PSOs. We used a male subject for the office environment and a female subject for the café. The gender assignments are random and are designed solely to provide variation across the environments. It does not convey or reinforce any potential gendered association biases. Figure 2 shows examples from both environments. The top row starts with the base café image containing only the face foreground PSO, followed by versions where medicine and then landmark are added as background PSOs. The bottom row starts with the base office image, followed by versions where the tattoo and then an additional face (a.) are added as background PSOs. The prompts we used for image generation are provided in Appendix A-A.